# Structured Probabilistic Coding

### 21 Dec 2023 v1

@@ Abstract
This paper presents a new supervised representation learning framework, namely Structured Probabilistic Coding (SPC), to learn compact and informative representations from input related to the target task. SPC is an encoder-only probabilistic coding technology with a structured regularization from the target label space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, the hidden representation is encoded into a Gaussian distribution space, while maximizing the prior entropy of latent representations concerning label space. This technique can simultaneously perform information encoding and task prediction in one module to more fully utilize the effective information from input data, and use variational inference in the output space to reduce randomness and uncertainty. To better control the probability distribution in the latent space, a structured regularization is proposed to promote class-level uniformity in the latent space. With the regularization term, SPC can preserve the Gaussian distribution structure of latent code as well as better cover the hidden space with class uniformly. We conduct evaluations on 12 natural language understanding tasks. The results show that our SPC can effectively improve the performance of pre-trained language models for various classification and regression tasks. Experiments demonstrate that SPC can enhance the generalization capability, robustness to label noise, and clustering quality of output representations.
@@ Introduction
Probabilistic embedding (Vilnis and McCallum 2015) is a flexible representation learning technology whose goal is to learn the underlying probability distribution of data. In contrast to deterministic embedding (Pereyra et al. 2017; Miyato, Dai, and Goodfellow 2017; Gunel et al. 2021), which maps each data to a fixed vector representation, probabilistic embedding embraces the notion of learning a probability distribution, mapping each data point to a distribution. These probabilistic embedding approaches can better describe the uncertainty and complexity of data, handle redundant information, and provide better discriminative representations. Probabilistic embedding has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Hu et al. 2022a).
@@
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are or can be built upon the information bottleneck (IB) principle (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that maximally preserves information about the output, striking a balance between compression and prediction. These IB-based methods typically involve two parametric modules, i.e., an encoder and a decoder. The encoder maps the input data to a probabilistic distribution in the latent space, and the decoder maps the probabilistic distribution to class probabilities for classification, predicted output values for regression, and generated samples for generative tasks. The encoder and decoder are trained jointly to minimize a task-specific loss function and maximize the mutual information between the input data and the latent space.
@@
However, under the encoder-decoder architecture, the process of mapping input data to probability distributions by the encoder may lose some task-related information which is essential for decoder. This is because probability distributions inherently contain randomness and uncertainty, which may be irrelevant to the task and interfere with the model’s learning of task-related information. To avoid this, we propose an encoder-only embedding technology, probabilistic coding, that combines probabilistic encoding and task prediction into one module. By using variational inference in the output space, we can better control and utilize randomness and uncertainty of data. This helps improve the model’s generalization performance, especially when facing incomplete or noisy labels. Combining probabilistic encoding and task prediction into one module can promote to learn more compact representations. These representations can better capture the underlying structure of data, thereby improving task prediction performance.
@@
Besides, although probabilistic embedding methods can capture data uncertainty and complexity, they are restricted to limited or biased data, which cannot fully represent the true distribution of the target task. Therefore, in the process of mapping input data to probability distributions in the latent space by the encoder, some task-related important information may be missing to some extent. The insufficient information may lead to limited task performance on new data and poor model generalization. To improve task prediction ability of the latent representations, we leverage the structured information of target task space to constrain the learning process of the latent space’s probability distribution. Under the framework of probabilistic coding, the structured regularization of latent space can help the model learn more informative representations related to the target task, thereby improving the model’s prediction accuracy on new data.
@@
In this paper, we present a new supervised representation learning framework, structured probabilistic coding (SPC), an encoder-only probabilistic coding technology with a structured regularization from the target label space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, we use variational approximation to encode hidden representation vectors into Gaussian distribution spaces while maximizing the prior information entropy of latent representations concerning label space. This technique can simultaneously perform information encoding and task prediction in one module to more fully utilize the effective information of input data, and use variational inference in the output space to reduce randomness and uncertainty. The method employs probabilistic encoding technique combined with labels to fit the true distribution of data well and encode the semantic structure of label spaces. At the same time, introducing the structure information of target task space during the probabilistic coding process to constrain the latent space’s probability distribution can make the probability distribution of latent space better reflect task-related information, which helps improve task prediction performance. The structured regularization promotes class-level uniformity in the latent space under the multivariate Gaussian distribution. With the regularization term, SPC can maintain the Gaussian structure of the neighborhood in the input space while achieving the best possible coverage of the hidden space with uniformity across classes.
@@
We conduct experiments on 12 natural language understanding tasks, including 10 classification tasks such as emoji prediction, social emotion detection, hate speech detection, irony detection, offensive language detection, sentiment analysis, and stance detection, as well as 2 regression tasks including semantic similarity prediction and plausible clarifications ranking. The results demonstrate that our SPC method can effectively improve the performance of pre-trained language models for classification and regression tasks. For instance, with the RoBERTa backbone, SPC can improve average performance by +4.0% and +1.5% for classification and regression tasks compared to CE. Our SPC framework consistently achieves the best average performance compared to other methods, including traditional deterministic methods (i.e., CE/MSE, CE/MSE+CP, CE+AT, and CE+SCL) and probabilistic methods (i.e. VIB, MINE-IB, and MEIB) under different backbone models such as BERT and RoBERTa. Extensive experiments show that SPC can enhance the pre-trained language models’ generalization capability including out-of-distribution and low-resource scenarios, robustness to label noise, and clustering quality of output representations.
@@
The main contributions are as follows: 1) We propose a novel encoder-only probabilistic coding method that combines probabilistic encoding and task prediction into one module. 2) We design a new structured regularization term to promote class-level uniformity in the latent space for better task prediction ability of probabilistic embedding. 3) We propose a supervised representation learning framework, namely SPC, to learn compact and informative representations from input related to the target task. 4) Experiments on 12 benchmarks show that SPC gains state-of-the-art performance on classification and regression tasks. Extensive experiments demonstrate that SPC can enhance the pre-trained language models’ generalization capability, robustness to label noise, and clustering quality of output representations. [^1] The code is available at https://github.com/zerohd4869/SPC
@@
The probabilistic coding combines probabilistic encoding and task prediction into one module. Different from existing probabilistic embeddings applying encoder-decoder architecture, our encoder-only paradigm can effectively maintain task-related features and avoid negative effects caused by the randomness and uncertainty of probabilistic encoding.
@@
Following Alemi et al. (2017), we employ a variational approximation to encode the input data into a Gaussian distribution space. Specifically, under the assumption that $p(t|x,y)=p(t|x)$, corresponding to the Markov chain $Y\rightarrow X\rightarrow T$, we aim to minimize the mutual information between the input $X$ and the latent representation $T$, i.e., $T\in\mathbb{R}^{|\mathcal{Y}|}$, as well as minimize the Kullback-Leibler (KL) divergence between the true posterior distribution $p(t|x)$ and the approximated distribution $r(t)$. The objective of probabilistic coding can be:
$\begin{split}\mathcal{L}_{PC}=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t)),\end{split}$
where $q(y|t)$ is a non-parametric operation, i.e., argmax function. $r(t)$ is an estimate of the prior $p(t)$ of $t$. $p_{\theta}(t|x)$ is the probability distribution for $x$, which is learned by the probabilistic encoder $\theta$. $\beta>0$ is a trade-off hyperparameter which controls the trade-off between the sufficiency of $t$ for predicting $y$, and the minimality of the representation $t$ from $x$. Here, $p_{\theta}(t|x)$ is typically set to the unit Gaussian $\mathcal{N}(0,I)$, i.e., $p_{\theta}(t|x)=\mathcal{N}(t|{\mu}(x),\Sigma(x))$, where $\mu$ and $\Sigma$ represent the mean and diagonal covariance, which are predicted by MLP with a single hidden layer. In probabilistic encoding, the re-parameterization trick (Kingma and Welling 2014) is used to make the model take gradients as usual.
@@
In existing IB-based methods (Alemi et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) with an encoder-decoder architecture, their decoder can be a parametric approximation $q_{\phi}$ of $p(y|t)$. That is, the compressed representation $t$ can be sampled from the distribution $p_{\theta}(t|x)$, meaning that a specific pattern of noise is added to the input of $q_{\phi}(y|t)$. This noise would decrease the information conveyed by $t$ and enable them to lose some task-related information, which is essential for the learning of decoder $\phi$. Different from them, our probabilistic coding applies a non-parametric operation to predict, and combines probabilistic encoding and task prediction into one encoder module with the network $\theta$. It can effectively avoid negative effects caused by the randomness and uncertainty of probabilistic encoding.
@@
As mentioned above, the Markov assumption restricts that the representation $T$ cannot depend directly on the target task $Y$. This means that the learning of $t$ does not fully utilize information of label space. Accordingly, the learned representation cannot sufficiently represent the true distribution of the target task, leading to poor generalization ability when learning from the limited or biased data. Therefore, we design a new structured regularization to explore underlying patterns of the label space.
Specifically, we add an additional term about the latent distribution to the objective function that maximizes the prior entropy of $T$ on the label space:
@@
In the implementation, we utilize each sampled batch data to estimate $H(T)$. The Jensen’s inequality is first applied to obtain a lower bound, i.e.,
$\begin{split}H(T)&=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{|\mathcal{Y}|}p_{i,j}\log p_{i,j}\\ &\geq-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{|\mathcal{Y}|}\frac{1}{B}\sum_{k=1}^{B}p_{k,j}\log\left(\frac{1}{B}\sum_{l=1}^{B}p_{l,j}\right),\\ \end{split}$
where $B$ is the batch size. $|\mathcal{Y}|$ is the number of classes. $p_{i,j}$ represents the probability of sample $i$ belonging to class $j$ in the latent space $T$, which is computed by the output of $p_{\theta}$, i.e., the encoder of probabilistic embedding. Then, the Monte Carlo method is used to estimate this lower bound:
$\begin{split}H(T)\geq\hat{H}(T)&\approx-\sum_{j=1}^{|\mathcal{Y}|}\left(\frac{1}{B}\sum_{k=1}^{B}p_{k,j}\right)\log\left(\frac{1}{B}\sum_{l=1}^{B}p_{l,j}\right)\\ &\approx-\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right)=\mathcal{L}_{b},\end{split}$
where $\overline{p_{\cdot,j}}=\frac{1}{N}\sum_{i=1}^{N}p_{i,j},j\in\{1,...,|\mathcal{Y}|\}$ represents the average predicted probability of the $j$-th target label variable. In this way, we can estimate $H(T)$ via computing the batch entropy $\mathcal{L}_{b}$. It measures the uncertainty or diversity of the predicted probability distribution over the label space. This allows us to introduce a structured regularization term that encourages uniformity across different classes in the latent space. After that, we can balance the learning of different features or labels, preventing the model from overemphasizing certain features or labels that are more prevalent in the training data but may not be representative of the true data distribution.
@@
$\begin{split}\mathcal{L}_{SPC}&=\mathcal{L}_{PC}-\gamma H(T)\gtrapprox\mathcal{L}_{PC}-\gamma\mathcal{L}_{b}\\ &=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t))\\ &\quad+\gamma\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right),\end{split}$
where $\gamma>0$ is a hyperparameter controlling the strength of the regularization. The first two terms combine probabilistic encoding and task prediction into one encoder module with the network $\theta$. The last regularization term promotes class-level uniformity in the latent space under the multivariate Gaussian distribution. Totally, the goal of SPC is to maintain the Gaussian structure of the neighborhood in the input space, as well as achieve the best possible coverage of the hidden space with uniformity across classes.
@@
Dataset Task # Label # Train # Val # Test Classification EmojiEval Emoji prediction 20 45,000 5,000 50,000 EmotionEval Social emotion detection 4 3,257 374 1,421 HatEval Hate speech detection 2 9,000 1,000 2,970 IronyEval Irony detection 2 2,862 955 784 OffensEval Offensive language detection 2 11,916 1,324 860 SentiEval Sentiment analysis 3 45,389 2,000 11,906 StanceEval Stance detection 3 2,620 294 1,249 ISEAR Emotion reaction prediction 7 3,066 767 3,833 MELD Conversational emotion recognition 7 9,989 1,109 2,610 GoEmotions Fine-grained emotion detection 28 36,308 4,548 4,591 Regression STS-B Semantic similarity prediction - 7,000 1,500 1,400 CLAIRE Plausible clarification ranking - 19,975 2,500 2,500
Table 1: The statistics of all datasets.
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. ${}^{\dagger}$ 29.30 64.70 36.70 61.70 52.30 62.90 67.30 - - - - ${}^{\dagger}$ 25.80 65.20 50.60 63.10 73.40 62.90 65.40 - - - - ${}^{\dagger}$ 24.70 66.00 52.60 62.80 71.70 58.30 59.40 - - - - GPT-3.5-turbo 6.34$\pm$0.01 73.23$\pm$0.18 48.30$\pm$0.11 66.81$\pm$3.26 63.71$\pm$0.13 40.40$\pm$3.13 39.45$\pm$0.10 67.22$\pm$0.09 41.46$\pm$0.11 25.21$\pm$0.08 47.21 BERT CE 22.30$\pm$0.60 76.05$\pm$1.41 44.67$\pm$1.78 59.38$\pm$3.01 80.16$\pm$1.26 70.54$\pm$0.44 65.21$\pm$0.71 67.17$\pm$0.78 39.80$\pm$0.84 46.29$\pm$0.79 57.16 CE+CP 21.91$\pm$0.71 76.28$\pm$1.20 45.97$\pm$2.93 64.06$\pm$2.41 78.99$\pm$1.57 70.68$\pm$0.31 65.83$\pm$0.39 67.20$\pm$0.95 39.54$\pm$1.61 46.39$\pm$0.63 57.69 CE+AT 22.93$\pm$0.70 75.08$\pm$1.23 46.30$\pm$3.61 64.23$\pm$2.04 79.68$\pm$1.59 70.55$\pm$0.57 66.46$\pm$1.13 65.70$\pm$0.69 39.84$\pm$0.38 47.37$\pm$0.54 57.81 CE+SCL 21.72$\pm$0.51 75.43$\pm$1.37 45.86$\pm$1.15 65.39$\pm$2.46 80.20$\pm$0.56 70.70$\pm$0.79 65.34$\pm$0.60 67.54$\pm$0.64 40.00$\pm$1.96 46.50$\pm$0.46 57.87 VIB 21.31$\pm$0.62 77.37$\pm$0.71 45.99$\pm$1.93 63.82$\pm$1.00 80.37$\pm$1.11 70.39$\pm$0.31 65.43$\pm$0.60 67.24$\pm$0.57 38.52$\pm$0.51 45.89$\pm$1.10 57.63 MINE-IB 21.29$\pm$0.31 76.60$\pm$0.41 47.64$\pm$2.11 65.86$\pm$2.57 78.67$\pm$2.28 69.85$\pm$0.54 65.35$\pm$0.88 67.62$\pm$0.40 41.23$\pm$0.67 46.87$\pm$0.42 58.10 MEIB 21.87$\pm$0.73 76.70$\pm$0.82 48.27$\pm$1.72 65.87$\pm$2.14 80.49$\pm$0.81 70.55$\pm$0.57 65.59$\pm$1.58 67.44$\pm$0.50 39.30$\pm$0.61 46.26$\pm$0.81 58.23 SPC (Ours) 24.19$\pm$1.55 77.15$\pm$0.73 57.48$\pm$2.99 65.85$\pm$1.07 80.65$\pm$0.78 70.74$\pm$0.12 67.17$\pm$1.08 68.94$\pm$0.35 42.68$\pm$0.94 47.62$\pm$1.38 60.25 RoBERTa CE 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43 CE+CP 31.12$\pm$0.84 77.54$\pm$0.70 48.59$\pm$3.28 58.75$\pm$6.19 79.50$\pm$0.98 72.82$\pm$0.29 66.89$\pm$1.67 70.58$\pm$0.71 40.74$\pm$0.89 47.98$\pm$0.65 59.45 CE+AT 32.00$\pm$0.93 77.30$\pm$1.07 44.71$\pm$4.76 60.17$\pm$3.17 79.81$\pm$1.11 72.51$\pm$0.44 67.81$\pm$0.95 70.97$\pm$0.68 40.10$\pm$0.60 47.89$\pm$1.21 59.33 CE+SCL 31.09$\pm$1.85 76.98$\pm$2.02 49.51$\pm$2.86 60.71$\pm$4.23 80.39$\pm$0.83 73.16$\pm$0.44 66.73$\pm$1.54 70.26$\pm$0.45 40.64$\pm$1.02 47.87$\pm$0.86 59.72 VIB 29.71$\pm$0.79 77.99$\pm$0.86 49.39$\pm$3.08 59.93$\pm$4.57 79.63$\pm$0.66 72.81$\pm$0.39 68.40$\pm$0.52 70.74$\pm$0.44 38.94$\pm$0.55 46.23$\pm$0.18 59.38 MINE-IB 31.70$\pm$0.45 78.79$\pm$0.58 46.39$\pm$2.82 57.39$\pm$8.27 79.76$\pm$0.67 72.85$\pm$0.56 67.27$\pm$1.00 70.15$\pm$0.58 41.80$\pm$2.14 48.88$\pm$1.04 59.50 MEIB 29.94$\pm$1.30 78.73$\pm$0.90 49.34$\pm$2.42 60.54$\pm$2.70 79.68$\pm$0.98 72.78$\pm$0.29 67.89$\pm$1.70 70.86$\pm$0.61 39.00$\pm$0.37 47.18$\pm$1.15 59.59 SPC (Ours) $*{}^{*}$$\pm$0.48 79.01$\pm$0.61 $*{}^{*}$$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 $*{}^{*}$$\pm$0.63 $*{}^{*}$$\pm$0.59 $*{}^{*}$$\pm$0.29 $*{}^{*}$$\pm$0.60 62.47
Table 2: Classification evaluation (%) on 10 benchmark datasets. ${}^{\dagger}$ means the results are from Barbieri et al. (2020). For other methods, we run five random seeds and report the average result on test sets. BERT and RoBERTa are the model backbones for deep representation learning technologies. Best results for each dataset are highlighted in bold. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
Methods STS-B CLAIRE Avg. Spearman Pearson Spearman Pearson MSE 88.33$\pm$0.32 88.80$\pm$0.36 47.10$\pm$3.36 48.06$\pm$3.43 68.07 MSE+CP 88.45$\pm$0.43 89.07$\pm$0.32 48.71$\pm$1.71 49.51$\pm$1.59 68.93 VIB 88.45$\pm$0.50 89.01$\pm$0.40 48.11$\pm$1.46 49.07$\pm$1.50 68.66 MEIB 88.61$\pm$0.14 89.13$\pm$0.17 48.97$\pm$1.86 47.87$\pm$1.83 68.64 SPC (Ours) 88.89$\pm$0.30 89.47$\pm$0.31 $*{}^{*}$$\pm$1.54 $*{}^{*}$$\pm$1.58 69.57
Table 3: Regression evaluation (%) on 2 datasets with RoBERTa backbone. For each method, we run five random seeds and report the average result on test sets. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
Methods EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions Avg. SPC 32.54$\pm$0.48 79.01$\pm$0.61 59.80$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 69.02$\pm$0.63 71.01$\pm$0.59 43.99$\pm$0.29 50.04$\pm$0.60 62.47   - w/o Structured 30.98$\pm$0.89 78.60$\pm$0.54 56.64$\pm$8.75 62.12$\pm$6.97 79.16$\pm$1.28 72.23$\pm$0.77 68.90$\pm$0.60 70.79$\pm$0.22 43.75$\pm$0.67 46.61$\pm$1.26 60.98   - w/o Structured - w/o Probabilistic 30.25$\pm$1.32 77.41$\pm$1.33 45.49$\pm$4.70 57.99$\pm$4.96 78.74$\pm$2.20 71.80$\pm$0.93 66.78$\pm$1.34 70.00$\pm$0.45 39.23$\pm$0.41 46.64$\pm$1.15 58.43
Table 4: Ablation results (%) on classification tasks. We experiment with RoBERTa backbone.

### 23 Feb 2024 v4

@@ Abstract
This paper presents a new supervised representation learning framework, namely structured probabilistic coding (SPC), to learn compact and informative representations from input related to the target task. SPC is an encoder-only probabilistic coding technology with a structured regularization from the target space. It can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, our probabilistic coding simultaneously performs information encoding and task prediction in one module to more fully utilize the effective information from input data. It uses variational inference in the output space to reduce randomness and uncertainty. Besides, to better control the learning process of probabilistic representations, a structured regularization is proposed to promote uniformity across classes in the latent space. With the regularization term, SPC can preserve the Gaussian structure of the latent code and achieve better coverage of the hidden space with class uniformly. Experimental results on 12 natural language understanding tasks demonstrate that our SPC effectively improves the performance of pre-trained language models for classification and regression. Extensive experiments show that SPC can enhance the generalization capability, robustness to label noise, and clustering quality of output representations.
@@ Introduction
Probabilistic embedding (Vilnis and McCallum 2015) is a flexible representation learning technology whose goal is to learn the underlying probability distribution of data. In contrast to deterministic embedding (Pereyra et al. 2017; Miyato, Dai, and Goodfellow 2017; Gunel et al. 2021), which maps each data to a fixed vector representation, probabilistic embedding embraces the notion of learning a probability distribution, mapping each data point to a distribution. These probabilistic embedding approaches can better describe the uncertainty and complexity of data, handle redundant information, and provide better discriminative representations. Probabilistic embedding has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Hu et al. 2022).
@@
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are (or can be) built upon the information bottleneck (IB) principle (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that preserves as much as possible information about the target task, striking a balance between compression and prediction. These IB-based methods typically involve two parametric modules, i.e., an encoder and a decoder (Goldfeld and Polyanskiy 2020). Usually, the encoder maps the input to a probabilistic distribution in the latent space, and the decoder maps the probabilistic distribution to the output representations in the target task space.
@@
However, under the encoder-decoder architecture, the process of mapping input data to probability distributions by the encoder may lose some task-related information, which is essential for the decoder during the learning process. This is because probability distributions inherently contain randomness and uncertainty, which may be irrelevant to the task and interfere with the task prediction process of the decoder. To avoid this, we propose an encoder-only embedding technology, probabilistic coding, that integrates probabilistic encoding and task prediction into one module. By using variational inference in the output space, we can better control and utilize randomness and uncertainty of data. The learned compact representations can fully capture the underlying structure of data, and preserve the effective information from input related to target task. This helps improve model generalization performance, especially when facing limited data or noisy labels.
@@
Besides, although probabilistic embedding methods can capture data uncertainty and complexity, they are restricted to limited or biased data, which cannot fully represent the true distribution of the target task. In the process of mapping input data to probability distributions in the latent space by the encoder, some task-related important information may be missing to some extent. The insufficient information lead to poor task performance on new data and inadequate model generalization. To improve task prediction ability of the latent representations, we leverage the structured information of target task space to constrain the learning process of the probability distribution in latent space. Under the framework of probabilistic coding, the structured regularization of latent space can help the model learn more informative representations related to the target task, thereby improving the model’s prediction accuracy on new data.
@@
In this paper, we present a new supervised representation learning framework, structured probabilistic coding (SPC), an encoder-only probabilistic coding technology with a structured regularization from the target label space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Specifically, the probabilistic coding technique performs variational approximation to encode the input into stochastic output representations under Gaussian distribution spaces, while minimizing the conditional entropy of the target label given the representations. Besides, the structure information of target task space is introduced to constrain the probability distribution of latent space. The structured regularization encourages class-level uniformity within the latent space under the multivariate Gaussian distribution, making the distribution better reflect task-related information, which is beneficial for task prediction. Under the probabilistic coding framework with the regularization term, SPC can maintain the Gaussian structure of the latent code while achieving the best possible coverage of the hidden space with uniformity across classes.
@@
We conduct experiments on 12 natural language understanding tasks, including 10 classification tasks such as emoji prediction, hate speech detection, irony detection, offensive language detection, sentiment analysis, stance detection, emotion detection from different domains, as well as 2 regression tasks including semantic similarity prediction and plausible clarifications ranking. The results demonstrate that our SPC effectively improves the performance of pre-trained language models for classification and regression tasks. For instance, with the RoBERTa backbone, SPC improves average performance by +4.0% and +1.7% for classification and regression tasks compared to CE/MSE. Our SPC framework consistently achieves the best average performance compared to other methods, including deterministic and probabilistic approaches, under different backbone models such as BERT and RoBERTa. Extensive experiments show that SPC can enhance the model generalization capability including out-of-distribution and data-constrained scenarios, robustness to label noise, and clustering quality of output representations.
@@
The main contributions are as follows: 1) We propose an encoder-only probabilistic coding method that integrates probabilistic encoding and task prediction into one module. It maximally preserves the effective information from input related to target task. 2) We design a structured regularization term to promote class-level uniformity in the latent space for better task prediction ability of probabilistic embedding. 3) We present a supervised representation learning framework named SPC, to learn compact and informative representations from input related to the target task. It can enhance the generalization ability of pre-trained language models for better language understanding. 4) Experiments on 12 benchmarks show that SPC achieves state-of-the-art performance on classification and regression tasks. Extensive experiments reveal that SPC can enhance the generalization capability, robustness to label noise, and clustering quality of output representations. [^1] The code is available at https://github.com/zerohd4869/SPC
@@
The probabilistic coding technology integrates probabilistic encoding and task prediction into one module. Different from existing probabilistic embeddings applying encoder-decoder architecture, our encoder-only paradigm can effectively maintain task-related features and avoid negative effects caused by the randomness and uncertainty of probabilistic encoding.
@@
Under the Markov chain constraint $Y\rightarrow X\rightarrow T$, we have $p(t|x,y)=p(t|x)$. we aim to minimize the mutual information between the input $X$ and the latent representation $T$, as well as maximize the information between the representations $T$ and the target label $Y$. Specifically, we employ a variational approximation to encode each input $x$ into a Gaussian distribution representation $t$ in the output space $\mathcal{Y}$, i.e., $T\in\mathbb{R}^{|\mathcal{Y}|}$. Additionally, we maximize the lower bound of $I(T;Y)$ by estimating the conditional entropy of the target label $Y$ given the representations $T$. The objective of probabilistic coding can be:
$\begin{split}\mathcal{L}_{PC}=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t)),\end{split}$ (1)
where $q(y|t)$ is a non-parametric operation, e.g., softmax function for classification. $r(t)$ is an estimate of the prior $p(t)$ of $t$. $p_{\theta}(t|x)$ is a variational estimate of the posterior probability of $t$ and is learned by the stochastic encoder $\theta$. $KL(\cdot)$ denotes the analytic KL-divergence term, serving as the regularization that forces the posterior probability of $t$ to approximately converge to the prior $p(t)$. $\beta>0$ is a hyperparameter which controls the trade-off between the sufficiency of $t$ for predicting $y$ and the compression of $t$ from $x$.
Let the prior $p(t)$ be the isotropic Gaussian distribution. And let the variational approximate posterior $p_{\theta}(t|x)$ be a multivariate Gaussian with a diagonal covariance structure, i.e., $p_{\theta}(t|x)=\mathcal{N}(t;{\mu}(x),\Sigma(x))$, where $\mu$ and $\Sigma$ represent the mean and diagonal covariance. Both of their parameters are input-dependent and predicted by an MLP (a fully-connected neural network with a single hidden layer), respectively. As the sampling of $t$ is a stochastic process, we apply the re-parameterization trick (Kingma and Welling 2014) to ensure unbiased gradients for the model.
@@
In existing IB-based methods (Alemi et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) with an encoder-decoder architecture, their decoder can be a parametric approximation $q_{\phi}$ of $p(y|t)$. That is, the compressed representation $t$ can be sampled from the distribution $p_{\theta}(t|x)$, meaning that a specific pattern of noise is added to the input of $q_{\phi}(y|t)$. The noise could diminish the information conveyed by $t$ and potentially cause a loss of task-related information, which is crucial for the decoder $\phi$ during the learning process. Different from them, our probabilistic coding integrates probabilistic encoding and task prediction into one encoder module with the network $\theta$, and applies a non-parametric operation to obtain prediction outputs. It effectively avoids negative impacts caused by the randomness and uncertainty of probabilistic encoding.
@@
As mentioned above, the Markov assumption restricts that the representation $T$ cannot depend directly on the target task $Y$. This means that the learning of $t$ does not fully utilize information of label space. Accordingly, the learned representation cannot sufficiently represent the true distribution of the target task, leading to poor generalization ability when learning from the limited or biased data. To enhance the prediction ability of the latent representation, we design a new structured regularization related to the target task space. It can constrain the learning process of probability distributions in the latent space.
Specifically, we add an additional regularization term about the latent distribution to the objective function that maximizes the prior entropy of $T$ on the label space:
@@
and
$\begin{split}H(T)&=-\frac{1}{|\mathcal{X}|}\sum_{i=1}^{|\mathcal{X}|}\sum_{j=1}^{|\mathcal{Y}|}p_{i,j}\log p_{i,j}\\ \end{split}$ (3)
where $|\mathcal{X}|$ is the number of the expected data. $|\mathcal{Y}|$ is the number of classes. $p_{i,j}$ represents the probability of sample $i$ belonging to class $j$ in the latent space, which is computed by the output of $p_{\theta}$, i.e., the encoder of probabilistic embedding.
In the implementation, the Jensen’s inequality is applied to estimate the upper bound of $H(T)$. Given each sampled batch data, we have:
$\begin{split}H(T_{b})\leq\hat{H}(T_{b})&\approx-\sum_{j=1}^{|\mathcal{Y}|}\left(\frac{1}{B}\sum_{k=1}^{B}p_{k,j}\right)\log\left(\frac{1}{B}\sum_{l=1}^{B}p_{l,j}\right)\\ &\approx-\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right)=\mathcal{L}_{b},\end{split}$ (4)
where $B$ is the batch size. $\overline{p_{\cdot,j}}=\frac{1}{N}\sum_{i=1}^{N}p_{i,j},j\in\{1,...,|\mathcal{Y}|\}$ represents the average predicted probability of the $j$-th target label variable. The Monte Carlo method are used to estimate the expected value of $H(T)$ via sampling batch data and computing the batch entropy $\mathcal{L}_{b}$, which measures the uncertainty or diversity of the predicted probability distribution over the label space. The regularization term encourages uniformity across different classes in the latent space. This approach allows for a balanced learning process across different labels, preventing the model from excessively emphasizing certain prevalent features within the training data that might not accurately represent the true data distribution.
@@
$\begin{split}\mathcal{L}_{SPC}&=\mathcal{L}_{PC}-\gamma H(T)\gtrapprox\mathcal{L}_{PC}-\gamma\mathcal{L}_{b}\\ &=\mathbb{E}_{t\sim p_{\theta}(t|x)}[-\log q(y|t)]+\beta KL(p_{\theta}(t|x);r(t))\\ &\quad+\gamma\sum_{j=1}^{|\mathcal{Y}|}\overline{p_{\cdot,j}}\log\left(\overline{p_{\cdot,j}}\right),\end{split}$ (5)
where $\gamma>0$ is a hyperparameter controlling the strength of the regularization. The first two terms combine probabilistic encoding and task prediction into one encoder module with the network $\theta$. The last regularization term promotes class-level uniformity in the latent space under the multivariate Gaussian distribution. Totally, the goal of SPC is to maintain the Gaussian structure of the latent code, as well as achieve the best possible coverage of the hidden space with uniformity across classes.